Presentation is loading. Please wait.

Presentation is loading. Please wait.

R.Giachino / M.Albert Be/op 2 th April 2015 LHC Machine check out 1v1.

Similar presentations


Presentation on theme: "R.Giachino / M.Albert Be/op 2 th April 2015 LHC Machine check out 1v1."— Presentation transcript:

1 R.Giachino / M.Albert Be/op 2 th April 2015 LHC Machine check out 1v1

2 Powering Groups of Circuits

3 Remaining HWC tests

4 Wednesday ● Communication lost in XR 5 ● EPC piquet to check gateway for triplet L5 ● Extended soft start in MKI2 & 8 successfully Normal soft start in MKI2 & 8 with sequencer tasks successfully To be done: – Pulse the MKI2 & 8 in injection test mode. For this we need to be at 450 GeV, so we need to wait until the four sectors are at injection 4

5 UL55 : Overheating ERD1/56 bus bars 5/32 Démonter la partie des jeux de barres endommagés ; Meuler un jeu de barre pour qu’il retrouve sa planéité (pour une surface de contact pleine et franche) ; Façonner une pièce d’adaptation si nécessaire (découpe, pliage, perçage) ; Nettoyer les jeux de barres et le tableau ; Remonter les parties ayant été démontées et réparées ; Effectuer une campagne complète de serrage sur l’ensemble du tableau ; Effectuer une test d’isolement, puis mis en service. Plan to be finished late afternoon-evening. problem 400 V (ERD1/56 ) TE/EPC in UL 55 - A bus-bar in an electrical switch board that has melted partially around a connection bolt. La partie endommagée est dû à un échauffement sur un jeu de barres en cuivre : deux pièces n’étaient plus en contact, et seul la boulonnerie assurait la passage du courant ; Lors de la mise sous tension d’un départ TE/EPC aujourd’hui, la planéité entre deux barres n’étant plus assurée, le défaut s’est révélé ;

6 UL55 : Overheating ERD1/56 bus bars 6/32

7 UL55 : Overheating ERD1/56 bus bars 7/32

8 UL55 : Overheating ERD1/56 bus bars 8/32

9 RB.A34 &RB.A67 ● Training quench RB.A34 at A (6.55 TeV). Frist magnet that quenched A17.L4 Followed by: B17.L4, A17.L4, A16.L4 ● 23:13 RB.A67 quench during a trim to A for the BETS test. Cell C31.L7, Looks like a training quench – Zinur will analyze tomorrow morning ● RB.A34 quench 1 minutes after Flat top – B8L4 and C8L4 -> propagate to A9L A -> 6.55TeV

10 RB.A34 during quench at 3.5 KA 10

11 Wednesday ● RCO &RCD’s 34 OK ● QPS threshold changed on RQT13/RQ6.R7, tested OK ● RB.A78 re-qualified ● RB.A34 re-qualified after earth fault in progress ● PIC-BIC only 4/5 circuits missing (in IP4) ● RD34 and RMSI powered to nominal, no problem appeared ● Ted’s Downstream moving IN /OUT – We call Mathieu Donze. TED are not moving. – Fault TED Watching dog time 11 6 sectors ramped no trips

12 Wednesday ● LHC closed ● Check the EIS and MSI ok ● Check BIS inputs from access & ASB. OK ● SIS interlock surveying the access is not OK – to be fixed (but not critical) 12

13 TeV ● BETS tracking reached correctly 6.5 TeV 13 - Both LBDS successfully ramped up to 6.5 TeV (RB45, RB56, RB67, RB78, Q4, MSD, MKD and MKB) - BEM interlocks OK for RB45 UP, RB45 DW, RB56 UP, RB56 DW, RB67 UP and RB67 DW - BEM interlocks for RB78 not tested due to trip of sector BEI interlocks OK for RQ4.B1 DW, RQ4.B2 DW, RMSD.B1 DW and RMSD.B2 DW - BEI interlocks NOT_OK for RQ4.B1 UP, RQ4.B2 UP, RMSD.B1 UP and RMSD.B2 UP - BEI tables clamped at 6.5 TeV. New tables for an energy up to 6.6 TeV to be generated and downloaded. Detailed data available in TE-ABT elogbook.

14 Tentative plan 14

15 Access requests WHOWHEREWHENWHATDuration LBDSPoint 6Q4 MSD Bets table update1 hr. EN/ELPoint 5ERD Repair5-8 hrs.

16 Plan of the day Access 10:00 – 17:00 EN/EL intervention on ERD1/56 in point 5 Repeat tests BETS tests at 6.5 TeV Tentative to close Beam Interlock Loop Vacuum valves open Remaining PIC- BIS test SPS Beam on Extraction TED. Nominal sequence tests Precycle, ramp and squeeze all sectors RF -Collimator LHC & Experiments closed after access 16 https://indico.cern.ch/category/6386/

17 Tentative plan ● Friday Sector 34 circuits operational – (If no problems encountered) ● Beam loop closing. (tonight) ● BIS machine protection tests. ● LBDS –BIS machine protection tests ● BIS Interlock valves (Tonight or Friday) ● All Machine cycling all sectors together ● First beam not before Easter… ● Checking Experts availability during Easter week end!! 17

18 Wednesday Exception when starting MKISS on MKI-8 from application MKISS however worked => Expert informed

19 MKI Summary of MKI test: Extended softstart in MKI2 & 8 - OK Normal softstart in MKI2 & 8 with sequencer tasks - OK To be done: pulse MKI2 & 8 in injection test mode => need to be at 450 GeV, i.e. need 4 sectors at injection Switching ON kickers with nominal sequence throws exception => MKI settings are now NON MULTIPLEXED, so task has to be changed

20 RF ● RF trip >> Problem fixed during the morning. – ACSLine8B2 ACSLine6B2 ACSLine5B2 ● Task to switch Rf Low level OFF and ON. when switching ON, the RF low level of line 5 B2 stays OFF(EXTRN RF VETO) 20

21 Wednesday Testing circuits in S34, except RB,RQF,RQD, RD34: Set to injection level OK 8:23 all sectors set to injection level Sequencer tasks fixed: SET EVENT GROUP eLOGBOOK COMMUNICATION Extended MKISS on MKI-2

22 Energy ● Reference energy correct! as we have ramped 56 and 67 ● SIS energy tracking at 6.5 TeV for Q4 & MSD ● Reception of all energies of the BLMs by the SIS is OK ! 22

23 Tuesday RQ6.R4B1 replaced and working correctly now (tripped on Monday evening on VS STATE fault) Intervention on RCBV4.R6B2 – OK First training quench on RCD/O in S34 RCO.A34B1/2RCD.A34B1/2

24 Tuesday – cycling Ramp of 7 sector w/o RCSX3.R1 and RB.A78 Trip of RQ5.R1B1/2

25 Tuesday – cycling BLM sanity check OK BLM 6.5TeV look OK Sequencer to e-logbook link broken – issue reported Another ramp of 7 sectors in the evening: NO trips !

26 Collimators – mp tests GV repeated all MP tests with updated software functionality showing the difference between LVDT and motor => no issues found, all OK

27 Collimators – test of new sequences Various issues found: Check Beta* limits for TCTs fails Load collimator settings to injection failed for TCT in IP2/IP8 Load collimator position thresholds fails Check of collimator injection settings fails for TCDQ, thresholds and TCTs in IP2/8 Ramped all UP and DOWN, left them at PARKING

28 Feedbacks wrong energy: OFC uses 480 GeV instead of 6.5 TeV. => to be fixed ! pattern of CODs read back from the FGCs are correct, but due to the wrong E, kicks are off by a factor 13 ! Tests of orbit feedback: Kick differences in H/V plane between OFC and LSA – OK 2 outliers: MCBX.3L8.B1 and MCBCH.B5L8.B1 MCBX.3L8: sign error (in LSA: POSITIVE for NEGATIVE kick) MCBCH.B5L8.B1: not present in run1, but fixed for run2 - suspicion that a configuration is wrong in the OFC since the OFB sees it as '0‘

29 Last night – cycling RAMPDOWN – RAMP – MEDIUM Beta UNSQUEEZE 19/30m + FT 7 sectors - all OK ! setting PC-FGC event group from sequencer doesn't work

30 Remaining issues Vacuum valve MSI-TI2 (VVGS.29528) interlock SPS EXT1 master BIC to be tested extraction conditions established >> Markus - BIC LHC probe beam re-phasing issues, stability issues => often BQM interlocks Threshold increased, to be checked with beam BTVDD image orientation > on hold until LBDS experts decide on preference BI-ABT decision to be made >> Jan -BI Controls: CMW/RDA issue affecting several clients new libraries released CIBDS problem while arming LBDS > Bruno, ongoing LBDS-XPOC: Missing BI data: (BCTFR, BSRA, etc..) >> Nicolas M- BI

31 Test to be done BIS LOOP closed ● Vacuum valves test with BIS loop closed ● LBDS BETS with 45 finished ● Circulate and Dump

32 Monday – cycling & PGCs RQ6.R4B1 tripped EPC >> RQ6.R4B1 ok now. RCBV4.R6B2 available again RQD.A45 tripped while going to STANDBY fault: VS_RUN_TO, OK after reset RQD/F.A78: I_MEAS fault => EPC needs to check it

33 BPM LHC SEQ: BPMLHC calibration finished. Overall result: ERROR Failed devices: LHC.BPMSI.A4L6.B2 LHC.BPMSI.A4R6.B1 LHC.BPMSX.A4L6.B1 LHC.BPMSX.A4R6.B2

34 Electrical P5 There is a busbar in an electrical switch board that has melted partially around a connection bolt. A specialist from EPC will go in to have a look and estimate the amount of work to repair it. As it looks now the repair will not be finished today

35 LHC closure test Switched EIS elements ON and powered to nominal while the machine was closed for 20 min. Everything worked well ! Trip in the end, because LHC access key was moved to avoid activating RP VETOs

36 HWC tests Combined powering of RCO/D.A34B2 to I_NOM + I_DELTA - OK PLEP to +I_NOM and -I_NOM on: RQT13.R7B1 OK RQT13.R7B2 OK RQ6.R7.B1 OK RQ6.R7.B2 OK

37 Training Quench RB.A A / 6.55 TeV First magnet that quenched: A17.L4 Followed by: B17.L4, A17.L4, A16.L4 T max : 23.67K

38 BETS 6.5 TeV BETS tracking reached 6.5 TeV

39 BETS 6.5 TeV Both LBDS successfully ramped up to 6.5 TeV (RB45, RB56, RB67, RB78, Q4, MSD, MKD and MKB) BEM interlocks OK for RB45 UP, RB45 DW, RB56 UP, RB56 DW, RB67 UP and RB67 DW BEM interlocks for RB78 not tested due to trip of S67 BEI interlocks OK for RQ4.B1 DW, RQ4.B2 DW, RMSD.B1 DW and RMSD.B2 DW BEI interlocks NOT OK for: RQ4.B1 UP, RQ4.B2 UP, RMSD.B1 UP and RMSD.B2 UP BEI tables clamped at 6.5 TeV => New tables for energy up to 6.6 TeV to be generated and downloaded.

40 RB.A67 trip during BETS test RB.67 tripped during a trim to A for BETS test Quench: cell C31.L7, magnet C30L7 T up to 24.3K => estimate: 7 h to recover cryogenics conditions

41 Check list before beam ● Close all hidden doors ● TED’s are Blocked 41


Download ppt "R.Giachino / M.Albert Be/op 2 th April 2015 LHC Machine check out 1v1."

Similar presentations


Ads by Google